the radius of gyration of a hemi sphere of mass 'm'and radius 'r' 


about an axis parallel to diameter at a distance '3r/4'is given by [center of mass of hemisphere lies at a height '3r/8' from base]


Grade:11

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear  bhargav ,

 Since the center of mass of a solid hemisphere is 3/8 R above the circular base; so reducing it to a one particle system, and the moment of inertia of that one particle system is M(3/8 R)^2=M(9/64)R^2.

 Now moment of inertia of the particle about an axis parallel to the original one but at a distance of 3/4 R can be obtained using the parallel axis theorem

 New M.I. = MR^2(45/64)

 So Radius of gyration= R(45/64)^1/2 
 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD